3y - x = 4
2x + 2y = 24
how do i solve this with substitution?

Answer by mananth(16949)


Rewrite the equations
-1x+3y= 4 .............1

2y=24-2x---------2
/2
y=12-x
Plug the value of y in (1)
-x+3(12-x)=4
-x+36-3x=4
-x-3x=4-36
-4x=-32
/-4
x=8
Plug the value of x in (1)
-x+3y=4
-1*8+3y=4
-8+3y=4
3y=12
/3
y=4